OpenStudy (anonymous):

solve for x (2x -3)/4x = 2/5

OpenStudy (anonymous):

you have to wonder if this website is helping people learn math or completely hindering it

OpenStudy (anonymous):

x = 7.5

OpenStudy (angela210793):

cross multiply 5(2x-3)=4x*2 10x-15=8x 10x-8x=15 2x=15 x=15/2

OpenStudy (anonymous):

2x-3 = (2/5)*4x 2x =(8x/5)+3 2x = (8x+15)/5 2x*5 = 8x+15 10x-8x = 15 2x = 15 x = 15/2 x = 15/2